Subject: Ignore previous minutes; they were from last year!

From: strauss@astro.Princeton.EDU

Submitted: Fri, 12 Feb 1999 11:54:54 -0500 (EST)

Message number: 338 (previous: 337, next: 339 up: Index)

	Apache Point Observatory 3.5m User's Committee Meeting
****************February 8, 1999**********************

Attending: Michael Strauss, Rene Walterbos, Bruce Gillespie, Ed
Turner, Jeff Brown, Connie Rockosi
Not attending: Alan Uomoto, Chris Stubbs, Ed Kibblewhite. 

****************Status of Instruments************************    

The AOTF, or Acousto-optical tunable filter, produced by John Hillman,
of Goddard, has been a visitor instrument a number of times at APO.  A
description of the instrument was posted by Nancy Chanover on
apo35-general last week.  It allows narrow-band imaging in the
near-IR.  It is available for general use this quarter; this is our
first time we've formally traded telescope time for the use of an
external instrument.  The instrument is complicated enough that it
requires having one of the experts around to make it work; contact
Nancy if you're interested in using it.  It will be available from
late April through late June this time, but may come back at various
times in the future (especially if there seems to be substantial
community interest). 

Negotiations are in progress for another IR imager from Sean Casey,
and for a Fabry-Perot from Bruce Woodgate. 

Gillespie: Dale Sandford at Yerkes is willing to come out during a dark
run soon, to do a "50,000 mile tuneup" of the GRIM electronics, to fix
some of the recent flakiness that has been seen recently.  Note that
this will not give the instrument any enhanced capabilities, but
should keep it alive for another year or so.

Stubbs and Magnier's NSF proposal for the DIS chip upgrade (better
readnoise, throughput, sampling, spectral coverage) was approved.
Hopefully, the upgrade will happen in calendar year 1999.  Perhaps we
should think about whether we could fix the DIS UV throughput problem
at the same time.  This is a conversation to have with Jim Gunn and
Mike Carr.

Status of echelle: Rockosi: People are quite happy; things look
stable: wavelength calibration in particular is stable.  The alignment
of the optics is good (2 pixel FWHM lines; in particular, the problem
seen in the Fall of asymmetric lines seems to have been fixed by
improved collimation of the optics); detector is doing well.  The slit
viewer is working within REMARK.  We're happy with the pointing model,
as well as the instrument block.  No autoguiding capability yet.  The
slit viewer is not as sensitive as we would like; at 15 mag, the
object disappears completely within the slit (making it a bit hard to
guide).  Observing fainter objects will therefore be difficult.  There
is a light leak that we are trying to plug. {Editor's Note: Connie
Rockosi has since diagnosed and fixed this problem.} Throughput is such that
we're roughly within a factor of 2 of what was hoped for (i.e., you
need to expose twice as long as the KPNO echelle for the same
signal-to-noise); it not yet clear where that factor of two is coming
from.  Could it be trouble with the tertiary?  Has the coating with
the grating degraded with time?  Read-noise is 6 e-, so one is
probably usually photon-limited.  The throughput in blue is not high;
but not bad either; you can work down at least to 3400 A.  Recoating
and/or washing of the telescope optics planned over the next few
months may buy back about 25% or so more throughput. 

  We need some discussion about operating this instrument remotely, as 
it is open for general use 2nd quarter!  Doug Duncan is working on
documentation (a user's manual); it is in progress. 

**********************News on the Mountain******************	     

Gillespie: Camron Hastings has given notice that he is leaving us, on
April 1.  We're interviewing people for open SDSS positions; many of
the applicants may be appropriate for the APO 3.5m observer, so it
may be possible to find a new observing specialist quickly. 

A "Community Meeting" was held on February 2 among those involved in
the various 3-year plan activities for the telescope.  We hope that this
is the first of a series of quarterly meetings of the APO staff,
engineers, etc. 
Minutes from this meeting are attached below as a Appendix 1 to the
present minutes.
Among the issues that came up:
  Karen Gloria has been looking at seeing as a function of wind
speed.  Below 15 mph, there is no correlation, but above this, the
seeing worsens by 0.3" for every 5 mph increase in windspeed.
Stiffening the secondary support may help this.
  Differential image motion monitor: Stubbs and a postdoc of his have
developed a system, which will come out to the site this week; useful
for 3.5m and SDSS. 
  A Shack-Hartman optical device will come later this month, along with
new secondary (see below).   It will be useful in collimation. 

  Steward is making real progress on the secondary; the worst points are
1.5 times spec.  They may finish polishing as soon as the end of this
month.  An e-mail from Steve Miller of the Mirror Lab is appended
as Appendix 2 below. 

  Gillespie: Any ideas of what to do with a used secondary?  Please
send him thoughts. 
Mechanicals for mounting the secondary will follow the secondary
itself; so sometime in March or April, we should be able to put it on. 

  Kibblewhite will start designing a second generation top end
(tip-tilt and other improvements); ready by the end of the year?  This
means that we could have both a spare secondary and spare top end in hand
by the end of the year. 

Tertiary rotation: Jeff Morgan is working hard; many of the parts are
in place.  Will be installed sometime in March. 

Telescope throughput: A witness sample for the primary was sent off to
QSP, which dis a proper reflectivity measurement.  It returned close
to nominal performance; 81-90% from UV through optical and near-IR.
This is after 2 years since the last aluminization.  So we're not
going to recoat the primary until at least Fall '99.   Work is ongoing
slowly thinking about getting our own aluminization tank. 

If you have anecdotes, etc., about the change in the throughput of the
system from your observing program over the past two years, please let
Gillespie know.  

Gillespie: Loomis says that NMSU is turning on the switch for vBNS
this week in Las Cruces.  Should be transparent to APO users; if all
goes well, it should not affect you at all (but things may be bumpy
right at first in practice).  The net result should be a great
improvement in throughput to APO users!

Everyone should be aware that the March 99 issue of Sky and Telescope
has 2 articles about 3.5m science! (one on high-redshift quasars, and
one on images of a low Galactic latitude galaxy). 

Last month's minutes are approved. 

March 8, 12 noon is next meeting. 

Appendix 1: 

This is a compacted summary of Mark Klaene's report of the 2/2/99 3.5-m
telecon  between APO staff and the UWash group--compiled by Bruce Gillespie

o General discusision of 3.5-m operations, including:

	- Identification of specific problem areas
		improve night-day communication (JD)
		relation between seeing vs windspeed (KG)
	- Feedback on split-shift observing spclst schedule
	- Status of Telescope Monitoring system  (Stubbs, Davis/Loomis)

    A few minor operational problems areas were identified:

	- need strap to assist in manually bringing telescope down in alt
	- automating the problem reporting system
	- better scattered light baffling is required
	- above 15 mph winds, we appear to degrade .3" in seeing every 5 mph

o Planned improvements projects for 1999:

	- Tertiary Rotation.  On schedule for a possible 10-day shutdown in
	  March.  An issue regarding eyelid operation was addressed by deciding
	  to implement the baseline system with manual operation of eyelids.

	- Telescope Monitoring system.  Expect 2ndary part of this system in
	  place by April possibly when the new secondary first installed.  Full
	  telescope monitor in work, expect az axis first, possibly this

	- Differential Image Motion Monitor. DIMM to be tested in Feb in
	  and out
	  of WSMR dome obs spec would like the data archived and available near
	  realtime. If nightly servicing amounts to 1 to 1 1/2 hours then Obs
	  spec see it as workable .

	- Interim mounting of new secondary (JonD). Holding band for coating in
	  fab, other support designs in work pending visit to SOML to confirm

	- Summer Shutdown. Wash primary, alt drive box rebuilt, az drive box
	  inspection, service PMSS pumps.

	- Front-end rebuild. Ed Kibblewhite to start design March. New features
	  to include: Stiffer supports, laser launcher, new radiation shield,
	  reduce heat load, tip-tilt actuators.

	- After tertiary rotation complete, UWash engineering to work on
	  improved baffling, echelle guiding, improved centroiding in general

o Also plan to devise means to improve and monitor collimation once
  Shack-Hartmann device arrives, in use about time new 2ndary installed

o Y2K, need to confirm compliance of computers, power, lpg, generator power,
  GPS/earth orientation data, Instruments, controllers.  Craig Loomis will plan
  for testing some systems during "March/3ry" shutdown. Jon Brinkman and Craig
  will look at FITS header info. DIS PC may be a problem, rest of the
  instruments appear OK.

o What to do with old secondary.spare as is, spare with refiguring, IR
secondary with top end.

o Outstanding items. Mirror cover automation, install new louvers for pit,
  counterweights for mirror covers.


Appendix 2: E-mail from Steve Miller of Steward Mirror Lab, Feb 6. 

We have completed the 6th iteration of figuring to the structure
function.  As of Thursday last, our figure looked like this:

34 @ .013
57 @ .025
81 @ .05
119@ .1
177 @ .2
205 @ .4

With the first number being the rms wavefront error in nm and the 
second, the spatial frequency in meters.  As you can see, we are about
1.5x out of spec on the middle frequencies.

Our first calculation of the structure function from a phase map (14
Jan) showed that all frequencies were out by a factor of 3 almost
uniformly.  For the first 4 iterations, the error dropped like a rock at
an average convergence rate of about14%.  We only got 2 runs in last
week since our optician, Bryan Smith, traveled back East for a wedding.

The last 2 iterations slowed a bit even taking a slightly backward step
at the 0.013 & 0.025 meter spatial frequencies.  We suspected a problem
with the lap calibration since the edge appeares to have gone down a bit
and are taking a day or two to do some tests.

I expect we will fix the problem in time to finish up this week or next.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 338 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to